OpenStudy (anonymous):

It takes Joan 5 imes longer than Jane to file the reports. Together they can file the reports in 10 minutes?

OpenStudy (anonymous):

you are missing some information in this question

OpenStudy (anonymous):

how long does it take each women to file the report

OpenStudy (gorv):

let jane take x minute jhon =5x

OpenStudy (gorv):

their per minute work 1/x and 1/5x

OpenStudy (gorv):

their total work in one minute = 1/10 1/x +1/5x=1/10

OpenStudy (gorv):

solve it